Two point charges Q1 = 44 C and Q2 = -5  C are separated by a distance of 17 cm. Find the electric potential energy of the system


1
Expert's answer
2022-03-02T10:17:11-0500

Potential energy is given by: PE = "\frac{k * Q1 * Q2}{r}"


where Q1, Q2 and r are the charge Q1, charge Q2 and distance between the two charges respectively.


Given, Q1 = 44C

Q2 = -5 C

r = 17 cm = 0.17 meter

and k = 8.99 * "10^9 \frac{Nm^2}{c^2}"


Therefore, PE = "\frac{8.99*10^9*44*(-5)}{0.17}"

On solving above, we get,

PE = -1.163 * "10^{13} J"